# Do Whitepaper Claims Predict Market Behavior? Evidence from Cryptocurrency Factor Analysis

### v2

@@ Abstract
Cryptocurrency projects articulate their value propositions through whitepapers, making foundational claims about functionality, use cases, and technical capabilities. This study investigates whether these narrative claims align with empirically observed market behavior. We construct a novel pipeline combining natural language processing (NLP) with tensor decomposition to compare three representational spaces: (1) a claims matrix derived from zero-shot classification of 38 whitepapers across 10 semantic categories using BART-MNLI, (2) a market statistics matrix capturing 7 financial metrics for 49 cryptocurrency assets over two years of hourly data (17,543 timestamps), and (3) latent factors extracted via CP tensor decomposition (rank 2, explaining 92.45% of variance). Using Procrustes rotation and Tucker’s congruence coefficient ($\phi$), we test alignment between narrative and market spaces across 37 common entities.
Results indicate weak alignment: claims–statistics ($\phi=0.246$, $p=0.339$), claims–factors ($\phi=0.058$, $p=0.751$), and statistics–factors ($\phi=0.174$, $p<0.001$). Critically, the statistics–factors significance ($p<0.001$) validates our methodology: since both matrices derive from the same market data, significant alignment confirms the Procrustes pipeline detects relationships when present, establishing that the narrative null result represents observed weak alignment rather than measurement failure. Inter-model validation using DeBERTa-v3 yields 32% exact agreement (Cohen’s $\kappa=0.14$) but 67% relaxed (top-3) agreement, indicating models capture similar semantic neighborhoods with different decision boundaries. Cross-sectional analysis reveals XMR, CRV, YFI, and SOL exhibit positive alignment contributions, while RPL, HBAR, AAVE, and SUSHI show the largest divergence.
We interpret these findings as consistent with weak alignment between whitepaper narratives and realized market factor structure. Limited statistical power ($n=37$) precludes distinguishing weak alignment from no alignment, but we can confidently reject strong alignment ($\phi\geq 0.70$). Implications for narrative economics, market efficiency, and investment analysis are discussed.
@@ 1 Introduction
Our findings reveal weak alignment across comparisons ($\phi<0.25$), with one notable exception: the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude, indicating systematic correspondence between market metrics and tensor-derived factors that narrative claims fail to capture. Specialized tokens (XMR, CRV, YFI, SOL) show positive alignment contributions, while DeFi infrastructure tokens (RPL, HBAR, AAVE, SUSHI) exhibit the largest narrative-market divergence. This result is robust to temporal variation, subsample perturbation, decomposition method (CP vs Tucker), and rank selection.
@@ 3.1 Market Data
We collect hourly OHLCV (open, high, low, close, volume) data from Binance for 49 cryptocurrency assets spanning January 1, 2023 to December 31, 2024, yielding 17,543 timestamps per asset. Asset selection follows liquidity and data availability criteria, including major cryptocurrencies (BTC, ETH) and a diverse set of DeFi, infrastructure, and utility tokens.
@@ 3.2 Whitepaper Corpus
We collect whitepapers for 38 assets where official foundational documents are publicly available: AAVE, ADA, ALGO, AR, ARB, ATOM, AVAX, BTC, COMP, DOT, ENS, ETH, FIL, GRT, ICP, LINK, MKR, NEAR, SC, SOL, STORJ, UNI, XMR, and ZEC. Documents include original whitepapers, consensus papers, protocol specifications, and technical documentation. PDF text is extracted using PyPDF2 with sentence-level tokenization; for assets without extractable PDFs, we use official documentation in markdown format. Table 2 summarizes corpus statistics for selected assets.
@@
+ 13 additional documents (see Appendix)
@@
Market data (Binance) 49 2-year hourly OHLCV
@@ 4.3.1 Zero-Shot Classification
We employ BART-large-MNLI (Lewis et al., 2020) for zero-shot text classification via the HuggingFace Transformers library. [^2] Model: facebook/bart-large-mnli. BART-large fine-tuned on Multi-Genre Natural Language Inference (MNLI; Williams et al. 2018) for natural language inference. Whitepapers are segmented into 500-word chunks (n=2,056n=2{,}056 across 24 entities); this chunk size balances computational efficiency with context preservation, following common practice in large-scale text classification. Each chunk is classified against ten domain-relevant categories. Zero-shot classification follows the entailment approach of Yin et al. (2019), constructing hypotheses of the form “This text is about [category]” for each candidate label.
@@ 4.3.3 Model Validation
Figure 1: Multi-method classification comparison across 10 semantic categories. Bar heights represent mean classification scores for BART-NLI (primary), sentence embeddings, and Ministral-3 LLM. Categories with high cross-method agreement (medium_of_exchange, smart_contracts, DeFi) show consistent relative rankings; categories with weak agreement (store_of_value, governance) exhibit larger inter-method variance.
@@ 4.7.4 Reproducibility
All stochastic procedures use fixed random seeds (seed = 42 for CP-ALS initialization, tensor operations, and permutation tests) to ensure reproducibility. Permutation tests use $B=1000$ iterations; bootstrap procedures use $B=1000$ resamples. Implementation uses Python 3.11+ with NumPy, SciPy, TensorLy (CP-ALS), scikit-learn (CCA, PLS), and Hugging Face Transformers (BART-MNLI). Code and data are available at [repository URL].
@@ 5.2 Claims Matrix
Figure 4: Claims matrix: Zero-shot classification scores across selected assets and 10 functional categories. Full corpus includes 38 assets; subset shown for readability.
@@ 5.3 Primary Alignment Tests
All three comparisons yield weak alignment ($\phi<0.25$). Notably, the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude. Statistical significance indicates the alignment is reliably above chance (non-random coupling), while weak magnitude ($\phi=0.174$, well below the 0.65 threshold) indicates this coupling is substantively negligible for practical inference. Importantly, both statistics and factors derive from the same underlying market data—summary statistics aggregate temporal behavior, while tensor factors capture latent structure—so the statistics–factors alignment serves as a calibration check: significant alignment ($p<0.001$) confirms our Procrustes pipeline successfully detects mathematical relationships when they exist, validating the methodology and establishing that the null result for claims represents a true negative rather than measurement failure. The key finding is that narrative claims show no such systematic relationship: claims-based comparisons yield both weak magnitude and non-significant p-values ($p=0.339$ and $p=0.751$), indicating that whatever market-data coupling exists does not extend to whitepaper content.
@@ 5.6 Temporal Stability
Alignment shows moderate variation throughout the sample period, ranging from $\phi=0.138$ (mid-2023, late-2024) to $\phi=0.200$ (late 2023). The expanded corpus ($n=37$) improves statistical power while revealing greater temporal heterogeneity.
@@ 6.7 Limitations
  • Single exchange (Binance) data may not represent broader market dynamics.
@@ 7 Conclusions
We investigated whether cryptocurrency whitepaper claims predict market behavior using a novel pipeline combining NLP claims extraction, tensor decomposition, and Procrustes alignment. Our analysis yields nuanced results: while narrative-market alignment remains weak ($\phi<0.25$), the statistics–factors comparison achieves statistical significance ($p<0.001$), suggesting market metrics systematically relate to tensor-derived factors even as narrative claims remain decoupled.

### v3

@@ Abstract
This study investigates whether cryptocurrency whitepaper narratives align with empirically observed market factor structure. We construct a pipeline combining zero-shot NLP classification of 38 whitepapers across 10 semantic categories with CP tensor decomposition of hourly market data (49 assets, 17,543 timestamps). Using Procrustes rotation and Tucker’s congruence coefficient ($\phi$), we find weak alignment between claims and market statistics ($\phi=0.246$, $p=0.339$) and between claims and latent factors ($\phi=0.058$, $p=0.751$). A methodological validation comparison—statistics versus factors, both derived from market data—achieves significance ($p<0.001$), confirming the pipeline detects real structure. The null result indicates whitepaper narratives do not meaningfully predict market factor structure, with implications for narrative economics and investor decision-making. Entity-level analysis reveals specialized tokens (XMR, CRV, YFI) show stronger narrative–market correspondence than broad infrastructure tokens.
@@ 1 Introduction
Our findings reveal weak alignment across comparisons ($\phi\leq 0.25$), with one notable exception: the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude, indicating systematic correspondence between market metrics and tensor-derived factors that narrative claims fail to capture. Specialized tokens (XMR, CRV, YFI, SOL) show positive alignment contributions, while DeFi infrastructure tokens (RPL, HBAR, AAVE, SUSHI) exhibit the largest narrative-market divergence. This result is robust to temporal variation, subsample perturbation, decomposition method (CP vs Tucker), and rank selection.
@@ 3.1 Market Data
We collect hourly OHLCV (open, high, low, close, volume) data via the CoinGecko API for 49 cryptocurrency assets spanning January 1, 2023 to December 31, 2024, yielding 17,543 timestamps per asset. Asset selection follows liquidity and data availability criteria, including major cryptocurrencies (BTC, ETH) and a diverse set of DeFi, infrastructure, and utility tokens.
@@ 3.2 Whitepaper Corpus
We collect whitepapers for 38 assets where official foundational documents are publicly available, including AAVE, ADA, ALGO, AR, ARB, ATOM, AVAX, BTC, COMP, DOT, ENS, ETH, FIL, GRT, ICP, LINK, MKR, NEAR, SC, SOL, STORJ, UNI, XMR, and ZEC (see Table 2 for the full asset list). Documents include original whitepapers, consensus papers, protocol specifications, and technical documentation. PDF text is extracted using PyPDF2 with sentence-level tokenization; for assets without extractable PDFs, we use official documentation in markdown format. Table 2 summarizes corpus statistics for selected assets.
@@
+ 27 additional documents (see Appendix)
@@
Market data (CoinGecko) 49 2-year hourly OHLCV
@@ 4.3.1 Zero-Shot Classification
We employ BART-large-MNLI (Lewis et al., 2020) for zero-shot text classification via the HuggingFace Transformers library. [^2] Model: facebook/bart-large-mnli. BART-large fine-tuned on Multi-Genre Natural Language Inference (MNLI; Williams et al. 2018) for natural language inference. Whitepapers are segmented into 500-word chunks (n=2,056n=2{,}056 across the initial 24-asset corpus); this chunk size balances computational efficiency with context preservation. The corpus was subsequently expanded to 38 assets as additional whitepapers were identified (see Table 2). Each chunk is classified against ten domain-relevant categories. Zero-shot classification follows the entailment approach of Yin et al. (2019), constructing hypotheses of the form “This text is about [category]” for each candidate label.
@@ 4.3.3 Model Validation
Figure 1: Multi-method classification comparison across 10 semantic categories. Bar heights represent mean classification scores for BART-NLI (primary), sentence embeddings, and Ministral-3 LLM. Categories with high cross-method agreement (medium_of_exchange, DeFi, privacy) show consistent relative rankings; categories with weak agreement (smart_contracts, store_of_value) exhibit larger inter-method variance.
@@ 4.7.4 Reproducibility
All stochastic procedures use fixed random seeds (seed = 42 for CP-ALS initialization, tensor operations, and permutation tests) to ensure reproducibility. Permutation tests use $B=1000$ iterations; bootstrap procedures use $B=1000$ resamples. Implementation uses Python 3.11+ with NumPy, SciPy, TensorLy (CP-ALS), scikit-learn (CCA, PLS), and Hugging Face Transformers (BART-MNLI). Code and data are available at https://github.com/studiofarzulla/tensor-defi.
@@ 5.2 Claims Matrix
Figure 4: Claims matrix: Zero-shot classification scores across selected assets and 10 functional categories. Full corpus of 38 assets shown.
@@ 5.3 Primary Alignment Tests
All three comparisons yield weak alignment ($\phi\leq 0.25$). Notably, the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude. Statistical significance indicates the alignment is reliably above chance (non-random coupling), while weak magnitude ($\phi=0.174$, well below the 0.65 threshold) indicates this coupling is substantively negligible for practical inference. Importantly, both statistics and factors derive from the same underlying market data—summary statistics aggregate temporal behavior, while tensor factors capture latent structure—so the statistics–factors alignment serves as a calibration check: significant alignment ($p<0.001$) confirms our Procrustes pipeline successfully detects mathematical relationships when they exist, validating the methodology and establishing that the null result for claims represents a true negative rather than measurement failure. The key finding is that narrative claims show no such systematic relationship: claims-based comparisons yield both weak magnitude and non-significant p-values ($p=0.339$ and $p=0.751$), indicating that whatever market-data coupling exists does not extend to whitepaper content.
@@ 5.4 Rank Sensitivity Analysis
Percentages rounded to nearest 0.1%; apparent duplicates reflect independent rounding of distinct values.
@@ 5.6 Temporal Stability
Alignment shows moderate variation throughout the sample period, ranging from $\phi=0.138$ (late 2024) to $\phi=0.200$ (late 2023), with all windows remaining in the weak-alignment range. The expanded corpus ($n=37$) improves statistical power while revealing greater temporal heterogeneity.
@@ 6.7 Limitations
  • Single data provider (CoinGecko) aggregated prices may not represent venue-specific microstructure dynamics.
@@ 7 Conclusions
We investigated whether cryptocurrency whitepaper claims predict market behavior using a novel pipeline combining NLP claims extraction, tensor decomposition, and Procrustes alignment. Our analysis yields nuanced results: while narrative-market alignment remains weak ($\phi\leq 0.25$), the statistics–factors comparison achieves statistical significance ($p<0.001$), suggesting market metrics systematically relate to tensor-derived factors even as narrative claims remain decoupled.
@@ 7.4 Final Remarks
Declarations
Funding
Supported by King’s College London resources. No external funding was received.
Conflict of Interest
The author declares no competing interests.
Ethics Approval
Not applicable.
Consent to Participate
Not applicable.
Consent for Publication
Not applicable.
Data Availability
Cryptocurrency market data obtained via CoinGecko API (publicly available). Whitepaper corpus collected from official project documentation (publicly available). Processed datasets and NLP classification outputs available at https://github.com/studiofarzulla/tensor-defi.
Code Availability
Full replication code available at https://github.com/studiofarzulla/tensor-defi under MIT License.
Author Contribution
Sole author.